OpenStudy (anonymous):

Given A(5,-8) and B(-6,2), find the points on segment AB that is three-fourths of the way from A to B.

OpenStudy (anonymous):

Find the midpoint which is (-1/2, -3) or (-0.5, -3). Then find the midpoint of this midpoint and point B which is (-13/4, -1/2) or (-3.25, -0.5)

OpenStudy (anonymous):

I don't understand why you did, what you did.

OpenStudy (zarkon):

I would do it like this ... .25(5,-8)+.75(-6,2)

OpenStudy (anonymous):

Why would you do it that way Zarkon?

OpenStudy (zarkon):

\[f(t)=(1-t)(5,-8)+t(-6,2)\] is the equation of the line segment (for \(0\le t\le 1\)) between A and B pick \(t=.75\) to go 3/4th the way to B

OpenStudy (anonymous):

Is that supposed to be "is" or "if" and I think I understand thank you